a pair of hikers travel 0.7km on a true bearing of 240 degrees and then 1.3km on a true bearing of 300 degree. How far west have they travelled from their starting point?
Answer by Alan3354(69443)

1191721
a pair of hikers travel 0.7km on a true bearing of 240 degrees and then 1.3km on a true bearing of 300 degree. How far west have they travelled from their starting point?
-------------------------
Starting at the Origin:
The first point (where they change headings) is the intersection of the line y = sqrt(3)*x and the circle x^2 + y^2 = 0.49
--> x^2 + 3x^2 = 0.49
x^2 = 0.49/4 = 0.1225
x = -0.35
y^2 = 0.49 - 0.1225 = 0.3675
y = ~ -0.6062
---> the point A at (-0.35,-0.6062)
==================
The end point is the intersection of the line thru point A with a slope of
-sqrt(3) and the circle (x+0.35)^2 + (y+0.6062)^2 = 1.3^2
The distance west is the x value at that point.
=================